### v1

@@ 4.1 Benchmarks
We formulate the Bijections and Functions benchmarks using two domains, $\Gamma$ and $\Delta$, as this formulation is known to help FOKC algorithms find efficient solutions [6]. To compare Gantry and ForcLift with FastWFOMC, which has no support for multiple domains, we set $|\Gamma|=|\Delta|$.
The three benchmarks cover a wide range of possibilities. The Friends & Smokers benchmark uses multiple predicates and only two variables without employing any constructs specific to only one of the (W)FOMC algorithms. The Functions benchmark, on the other hand, can still be handled by all the algorithms, but its formulation relies on algorithm-specific constructs. Namely, for ForcLift and Gantry, the sentence is written using three variables. Since FastWFOMC supports at most two variables, we rewrite the Equation 7 with two variables and a cardinality constraint (see Appendix A for details about translating the benchmarks to other first-order logics). Lastly, the Bijections benchmark is an example of a sentence that FastWFOMC can handle but ForcLift cannot.
4.2 Results
Figure 2: The runtime of the algorithms as a function of domain size. Note that both axes are on a logarithmic scale.
Figure 2 presents a summary of the experimental results. Only FastWFOMC and Gantry-BFS could handle the bijection-counting problem. For this benchmark, the largest domain sizes these algorithms could accommodate were $64$ and $4096$, respectively. On the other two benchmarks, ForcLift had the lowest runtime. However, since it can only handle model counts smaller than $2^{31}$, it only scales up to domain sizes of $16$ and $128$ for Friends & Smokers and Functions, respectively. FastWFOMC outperformed ForcLift in the case of Friends & Smokers, but not Functions, as it could handle domains of size $1024$ and $64$, respectively. Furthermore, both Gantry-BFS and Gantry-Greedy performed similarly on both benchmarks. Similarly to the Bijections benchmark, Gantry significantly outperformed the other two algorithms, scaling up to domains of size $8192$ and $67,108,864$, respectively.
[…]
Based on our experiments, which algorithm should one use in practice? If ForcLift can handle the sentence and the domain sizes are reasonably small, it is likely the fastest algorithm. In other situations, Gantry will likely be significantly more efficient than FastWFOMC regardless of domain size, provided both algorithms can handle the sentence.
5 Conclusion and Future Work
In this work, we have presented a scalable, automated FOKC-based approach to FOMC. Our algorithm involves completing the definitions of recursive functions and subsequently translating all function definitions into C++ code. Empirical results demonstrate that Gantry can scale to larger domain sizes than FastWFOMC while supporting a wider range of sentences than ForcLift. The ability to efficiently handle large domain sizes is particularly crucial in the weighted setting, as illustrated by the Friends & Smokers example, where the model captures complex social networks with probabilistic relationships. Without this scalability, these models would have limited practical value.
Future directions for research include conducting a comprehensive experimental comparison of FOMC algorithms to better understand their comparative performance across various sentences. The capabilities of Gantry could also be characterised theoretically, for example, by proving completeness for logic fragments liftable by other algorithms [11, 22, 26]. Additionally, the efficiency of FOMC algorithms can be further analysed using fine-grained complexity, which would provide more detailed insights into the computational demands of different sentences.
Appendix A The Three Logics of FOMC
Logic Sorts Constants Variables Quantifiers Additional atoms
$\mathsf{FO}$ one or more unlimited $\forall$, $\exists$ $x=y$
$\mathsf{UFO}^{2}+\mathsf{CC}$ one two $\forall$ $|P|=m$
FOMC commonly utilises three types of first-order logic: $\mathsf{FO}$, $\mathsf{C}^{2}$, and $\mathsf{UFO}^{2}+\mathsf{CC}$. Table 1 summarises the key differences among them. $\mathsf{FO}$ is the input format for ForcLift and its extensions Crane and Gantry. $\mathsf{C}^{2}$ is often used in the literature on FastWFOMC and related methods [11, 12]. (Note that no algorithm accepts $\mathsf{C}^{2}$ as input.) Finally, $\mathsf{UFO}^{2}+\mathsf{CC}$ is the input format supported by the most recent implementation of FastWFOMC [23]. All three logics are function-free, and domains are always assumed to be finite. As usual, we presuppose the unique name assumption, which states that two constants are equal if and only if they are the same constant [18].
In $\mathsf{FO}$, each term has a designated sort, and each predicate $P/n$ corresponds to a sequence of $n$ sorts. Each sort has its corresponding domain. These assignments to sorts are typically left implicit and follow from the quantifiers, e.g., $\forall x,y\in\Delta\text{. }P(x,y)$ implies that the variables $x$ and $y$ have the same sort. On the other hand, $\forall x\in\Delta\text{. }\forall y\in\Gamma\text{. }P(x,y)$ implies that $x$ and $y$ have different sorts, and it would be improper to write, for example, $\forall x\in\Delta\text{. }\forall y\in\Gamma\text{. }P(x,y)\lor x=y$. $\mathsf{FO}$ is also the only logic to support constants, sentences with more than two variables, and the equality predicate. While we do not explicitly refer to sorts in the paper, the many-sorted nature of $\mathsf{FO}$ is paramount to the algorithms presented therein.
{remark*}
Compared to $\mathsf{FO}$, $\mathsf{C}^{2}$ and $\mathsf{UFO}^{2}+\mathsf{CC}$ lack support for constants, the equality predicate, multiple domains, and sentences with more than two variables. The advantage that $\mathsf{C}^{2}$ brings over $\mathsf{FO}$ is the inclusion of counting quantifiers. That is, alongside $\forall$ and $\exists$, $\mathsf{C}^{2}$ supports $\exists^{=k}$, $\exists^{\leq k}$, and $\exists^{\geq k}$ for any positive integer $k$. For example, $\exists^{=1}x\text{. }\phi(x)$ means that there exists exactly one $x$ such that $\phi(x)$, and $\exists^{\leq 2}x\text{. }\phi(x)$ means that there exist at most two such $x$. $\mathsf{UFO}^{2}+\mathsf{CC}$, on the other hand, does not support any existential quantifiers but instead incorporates (equality) cardinality constraints. For example, $|P|=3$ constrains all models to have precisely three positive literals with the predicate $P$.

### v2

@@ 4.1 Benchmarks
We formulate the Bijections and Functions benchmarks using two domains, $\Gamma$ and $\Delta$, as this formulation is known to help FOKC algorithms find efficient solutions [6]. To compare Crane2 and ForcLift with FastWFOMC, which has no support for multiple domains, we set $|\Gamma|=|\Delta|$.
@@
The three benchmarks cover a wide range of possibilities. The Friends & Smokers benchmark uses multiple predicates and only two variables without employing any constructs specific to only one of the (W)FOMC algorithms. The Functions benchmark, on the other hand, can still be handled by all the algorithms, but its formulation relies on algorithm-specific constructs. Namely, for ForcLift and Crane2, the sentence is written using three variables. Since FastWFOMC supports at most two variables, we rewrite the Equation 7 with two variables and a cardinality constraint. Lastly, the Bijections benchmark is an example of a sentence that FastWFOMC can handle but ForcLift cannot.
4.2 Results
Figure 2: The (total) runtime of the algorithms as a function of domain size. Note that both axes are on a logarithmic scale.
Figure 2 presents a summary of the experimental results. Only FastWFOMC and Crane2-BFS could handle the bijection-counting problem. For this benchmark, the largest domain sizes these algorithms could accommodate were $64$ and $4096$, respectively. On the other two benchmarks, ForcLift had the lowest runtime. However, since it can only handle model counts smaller than $2^{31}$, it only scales up to domain sizes of $16$ and $128$ for Friends & Smokers and Functions, respectively. FastWFOMC outperformed ForcLift in the case of Friends & Smokers, but not Functions, as it could handle domains of size $1024$ and $64$, respectively. Furthermore, both Crane2-BFS and Crane2-Greedy performed similarly on both benchmarks. Similarly to the Bijections benchmark, Crane2 significantly outperformed the other two algorithms, scaling up to domains of size $8192$ and $67,108,864$, respectively.
Table 2: The compilation times of FOKC algorithms, sorted in ascending order.
Algorithm Benchmark Compilation time (s)
Functions 0.44
ForcLift Friends & Smokers 0.52
Friends & Smokers 2.20
Crane2-Greedy Functions 2.36
Friends & Smokers 2.23
Functions 2.53
Crane2-BFS Bijections 7.70
One might notice that the runtime of FastWFOMC and ForcLift is slightly higher for the smallest domain size. This peculiarity is the consequence of just-in-time (JIT) compilation. As Crane2 is only run once per benchmark, we include the JIT compilation time in its overall runtime across all domain sizes. Additionally, Table 2 summarises the compilation time of each FOKC algorithm on each suitable benchmark. While the compilation time of ForcLift is generally lower compared to Crane2, neither significantly affects the overall runtime.
Based on our experiments, which algorithm should one use in practice? If ForcLift can handle the sentence and the domain sizes are reasonably small, it is likely the fastest algorithm. In other situations, Crane2 will likely be significantly more efficient than FastWFOMC regardless of domain size, provided both algorithms can handle the sentence.
5 Conclusion and Future Work
In this work, we have presented a scalable, automated FOKC-based approach to FOMC. Our algorithm involves completing the definitions of recursive functions and subsequently translating all function definitions into C++ code. Empirical results demonstrate that Crane2 can scale to larger domain sizes than FastWFOMC while supporting a wider range of sentences than ForcLift. The ability to efficiently handle large domain sizes is particularly crucial in the weighted setting, as illustrated by the Friends & Smokers example, where the model captures complex social networks with probabilistic relationships. Without this scalability, these models would have limited practical value.
Future directions for research include conducting a comprehensive experimental comparison of FOMC algorithms to better understand their comparative performance across various sentences. The capabilities of Crane2 could also be characterised theoretically, for example, by proving completeness for logic fragments liftable by other algorithms [11, 22, 26]. Additionally, the efficiency of FOMC algorithms can be further analysed using fine-grained complexity, which would provide more detailed insights into the computational demands of different sentences.